# Open problems and perspectives on solving Friedrichs systems by Krylov approximation

### v1

@@ Abstract.
Partially supported by the Italian National Institute for Higher Mathematics INdAM and the Alexander von Humboldt Foundation, Bonn.
@@ 1. Introduction
This clearly encompasses a huge variety of concrete classes of inverse problems, including those where $A$ is a differential operator, and the one hand it is customary to study each such class developing ad hoc methods and tools. This falls within the realm of numerical analysis, and includes the issue of reproducing the ‘exact’ solution(s) $f$ to $Af=g$ as a limit of approximants $f_{n}\to f$ obtained by solving (easier) truncated problems $A_{n}f_{n}=g_{n}$, typically in an iterative way.
@@
Obviously, there is a natural interplay between the two perspectives above. For instance, the actual convergence $f_{n}\to f$ from solutions to suitably truncated inverse problems obtained in a concrete setting, as well as the convergence rate and the quantitative control (say, the norm) of the displacement $f-f_{n}$, are often established by means of abstract functional-analytic and operator theoretic methods. We mention, in this respect, excellent classical references such as [45, 50, 56, 40, 28, 24, 41, 46, 52, 33, 9, 54, 51], and also, to quickly grasp the flavour of the abstract point of view, [18, Sections 1.1 and 1.2, and Appendix A].
@@
As a matter of fact, the theoretical apparatus developed so far in the subject of Krylov solvability is sufficiently articulated to make it natural to specialise it for distinguished classes of inverse problems. The one we consider here is that type of inverse problems customarily referred to as Friedrichs systems.
@@
Starting from concrete, differential formulations of Friedrichs-type inverse problems it is possible to re-formulate them in abstract Hilbert space setting, so as to exploit general operator-theoretic methods and then apply the results to concrete versions [33, 37, 1, 3, 7, 29, 29, 30]. Section 3 revisits the basic aspects of this construction.
@@ 2.1. Krylov approximation methods as truncation algorithms
(2.3) $\mathcal{K}_{n}(A,g)\,:=\,\mathrm{span}\big{\{}g,Ag,A^{2}g,\dots,A^{n-1}g\big{\}}\,,\qquad\dim\mathcal{K}_{n}(A,g)=n\,,$
@@ 2.2. A glance at Krylov solvability for bounded inverse problems
Here Krylov reducibility indicates that $A$ is reduced (in the standard operator sense, e.g., [53, Section 1.4]) into two blocks with respect to the Hilbert space direct sum decomposition
@@
An evident sufficient condition for Krylov reducibility is self-adjointness: a bounded self-adjoint operator $A$ is always $\mathcal{K}(A,g)$-reduced for any $g\in\mathcal{H}$. More generally, bounded normal operators $A$ are $\mathcal{K}(A,g)$-reduced if and only if $A^{*}g\in\overline{\mathcal{K}(A,g)}$ [21, Proposition 2.1]. Thus, for unitary operators Krylov reducibility is equivalent to Krylov solvability [21, Remark 3.6].
@@
[21, Proposition 3.9] Let $A\in\mathcal{B}(\mathcal{H})$ and let $g\in\mathrm{ran}A$. If A is $\mathcal{K}(A,g)$-reduced, then there exists a Krylov solution to the problem $Af=g$. For example, if $f^{\circ}\in\mathcal{H}$ satisfies $Af^{\circ}=g$ and $P_{\mathcal{K}}$ is the orthogonal projection onto $\overline{\mathcal{K}(A,g)}$, then $f^{\circ}:=P_{\mathcal{K}}f$ is a Krylov solution
@@
[21, Proposition 3.4], [17, Proposition 2.4] Let $A\in\mathcal{B}(\mathcal{H})$ and let $g\in\mathrm{ran}A$. If $\mathcal{I}(A,g)=\{0\}$, then there exists a Krylov solution $f\in\overline{\mathcal{K}(A,g)}$ to the inverse linear problem $Af=g$. If, additionally, $A^{-1}\in\mathcal{B}(\mathcal{H})$ then $\mathcal{I}(A,g)=\{0\}$ if and only if the solution $f$ to $Af=g$ is in $\overline{\mathcal{K}(A,g)}$.
Observe that, obviously, the $\mathcal{K}(A,g)$-reducibility of an operator $A$ is just a special case of the triviality of the Krylov intersection $\mathcal{I}(A,g)$, which indicates that the latter subspace is the more informative concerning the possibility that an inverse problem be Krylov-solvable.
@@
[23, Propositions 3.4 and 3.9] Let $A\in\mathcal{B}(\mathcal{H})$ be a compact normal operator and let $g\in\mathcal{H}$. Then $A$ is $\mathcal{K}(A,g)$-reduced. If, additionally, $g\in\mathrm{ran}A$, then $Af=g$ has a unique and minimal norm Krylov solution.
@@ 2.3. Krylov solvability and unbounded self-adjoint operators
and a discussion of its properties and relevance for the Krylov solvability of $Af=g$ may be found in [17, Sections 5 and 6].
@@
The self-adjoint case exhibits in this respect distinguished properties. For one thing, as well known (see, e.g., [53, Lemma 7.13]), if $A$ is self-adjoint, then $\mathcal{D}^{b}(A)$ is dense in $\mathcal{H}$, and so too are, consequently, all the other classes above. Moreover, a noticeable conclusion can be drawn concerning the Krylov solvability of possibly unbounded self-adjoint inverse problems.
@@
[17, Theorems 4.1 and 4.2] Let $A$ be a self-adjoint or skew-adjoint operator on a Hilbert space $\mathcal{H}$, and let $g\in\mathrm{ran}A\cap\mathcal{D}^{qa}(A)$. Then there exists a unique Krylov solution to the inverse linear problem $Af=g$.
It is worth commenting that Theorem 2.7 relies on an ingenious and somewhat laborious analysis carried out in the framework of the convergence of the conjugate gradient algorithm on an infinite-dimensional Hilbert space. This was initially developed in the bounded case [48, 49] and recently re-designed in the unbounded setting [19, Theorem 2.4 and Corollary 2.5]. The full proof obtained in [19] in fact reveals that the sequence of approximants from the conjugate-gradients algorithm converges to the minimal norm solution of the inverse problem (2.1) when $A$ is positive self-adjoint and $g\in\mathcal{D}^{qa}(A)\cap\mathrm{ran}A$. In other words: the truncation scheme and truncated inverse problems (2.4) defined by setting the trial space $Q_{n}=P_{n}$ for $P_{n}$ the orthogonal projection onto $\mathcal{K}_{n}(A,g)$, produce a well-defined sequence of approximants $(f_{n})_{n\in\mathbb{N}}$ and $f_{n}\xrightarrow{n\to\infty}f^{\circ}$ in $\mathcal{H}$, where $f^{\circ}$ is the minimal norm solution to (2.1).
@@ 2.4. Krylov subspaces and solvability under perturbations
[20, Example 3.3] Let $R:\ell^{2}(\mathbb{Z})\to\ell^{2}(\mathbb{Z})$ be the right shift operator $e_{n}\mapsto e_{n+1}$ (where $e_{n}$ is the $n$-th canonical vector of $\ell^{2}(\mathbb{Z})$). $R$ is unitary with $R^{*}=R^{-1}=L$, the left-shift operator $e_{n}\mapsto e_{n-1}$. $R$ admits a dense set $\mathcal{C}\subset\ell^{2}(\mathbb{Z})$ of cyclic vectors (i.e., vectors $g\in\mathcal{H}$ such that $\mathcal{K}(R,g)$ is dense in $\ell^{2}(\mathbb{Z})$) and a dense subset $\mathcal{N}\subset\ell^{2}(\mathbb{Z})$ of all finite linear combinations of the canonical vectors. Clearly all vectors in $\mathcal{N}$ are non-cyclic for the operator $R$, and moreover the solution $f$ to $Rf=g$ for any $g\in\mathcal{N}$ does not belong to $\overline{\mathcal{K}(R,g)}$.
@@
In a perturbative perspective, it is relevant to compare the possibly “close” Krylov subspaces $\mathcal{K}(A,g)$ and $\mathcal{K}(\widetilde{A},\widetilde{g})$. A promising tool is the weak gap metric between subspaces of separable Hilbert space, inspired to the standard norm gap distance (namely, the Hausdorff distance) between subspaces (see, e.g., [44, Chapt. 4, §2]). Owing to the separability, and hence reflexivity, of $\mathcal{H}$, the Hilbert space weak topology restricted to the (norm-closed) unit ball $B_{\mathcal{H}}$ of $\mathcal{H}$ makes the ball a compact and complete metric space $(B_{\mathcal{H}},\rho_{w})$, and the metric $\rho_{w}$ is induced by a norm, namely $\rho_{w}(x,y)=\|x-y\|_{w}$ for some explicitly constructed norm in $B_{\mathcal{H}}$. This allows to define, for any pair of norm-closed linear subspaces $U$ and $V$ of $\mathcal{H}$,
@@
where $B_{U}=B_{\mathcal{H}}\cap U$ and $B_{V}=B_{\mathcal{H}}\cap V$. As a matter of fact [20, Lemmas 6.1 and 6.2], $\widehat{d}_{w}$ is indeed a metric for the norm-closed subspaces of $\mathcal{H}$, albeit in general it is not complete.
@@
of the $N$-dimensional truncated Krylov subspace $\mathcal{K}_{N}(A,g)$ to the (closure of the) full Krylov subspace $\mathcal{K}(A,g)$. Such quite natural inner Krylov approximability is actually false when the standard Hausdorff distance is used for an infinite-dimensional $\mathcal{H}$ [20, Lemma 7.1]. Another natural aspect of the inner Krylov approximability that manifests in the $\widehat{d}_{w}$-sense is the following [20, Lemma 7.6]:
@@
Other natural approximations, instead, do not hold in general in the $\widehat{d}_{w}$-sense, which suggests that further conditions need be added to the perturbation picture. For example [20, Lemma 7.3, Examples 7.4 and 7.5],
@@
[20, Proposition 7.7] Let $\mathcal{H}$ be a separable Hilbert space, and let the following be given:
@@ 3. Overview on abstract Friedrichs systems
They are currently being investigated in a variety of directions, including well-posedness [37, 1, 3, 7], homogeneisation [13, 12], representations of boundary conditions [1], connection with the classical theory [2, 3, 6, 4], applications to various initial or boundary value problems of elliptic, hyperbolic, and parabolic type [5, 14, 26, 34, 47], development of different numerical schemes [11, 10, 15, 34, 35, 36], classification of abstract realisations of Friedrichs systems [7, 29, 30, 31, 32], the Friedrichs system structure of contact interactions [29], and many others.
@@
Many other concrete differential problems may be cast in the form $Af=g$ with $A$ of the type (3.1)-(3.2) [1, 2, 6, 5, 14]. In addition, it is possible to reformulate all such concrete problems at an abstract Hilbert space setting: this includes the classification of the distinct operator realisations of $A$ (namely, the counterpart of boundary conditions at $\partial\Omega$), as well as the identification of certain intrinsic abstract conditions ensuring the bijectivity of $A$ and hence the solvability of $Af=g$ [37, 1, 3, 8, 29, 30, 31, 32].
@@
In the abstract setting (3.3) it can be shown ([37, Section 2.1], [3, Theorem 7]) that the graph norms associated with $A_{0}$ and $\widetilde{A}_{0}$ are equivalent, and as a consequence $A_{0}$ and $\widetilde{A}_{0}$ have a common (dense) domain
@@
(3.6) $A_{1}+\widetilde{A}_{1}\,=\,\overline{A_{0}+\widetilde{A}_{0}}\,.$
@@
In particular, a special status is attributed to those $V$’s that are closed in $(\mathcal{W}_{1},\|\cdot\|_{A_{1}})$ and correspondingly those realisations $A_{1}|_{V}$’s that are bijective as $(V,\|\cdot\|_{A_{1}})\to\mathcal{H}$ maps. Indeed, such operators are closed and densely defined in $\mathcal{H}$, are also $(V,\|\cdot\|_{A_{1}})\to\mathcal{H}$ bounded, and their inverse is $\mathcal{H}\to(V,\|\cdot\|_{A_{1}})$ bounded, implying that the inverse linear problem induced by each such $A_{1}|_{V}$ is well-posed. Incidentally, one can see that if $A:=A_{1}|_{V}$ is a closed bijective realisation of $A_{0}$, then its adjoint $A^{*}=(A_{1}|_{V})^{*}\equiv\widetilde{A}_{1}|_{\widetilde{V}}$, with $\widetilde{V}:=\mathrm{dom}\,(A_{1}|_{V})^{*}$, is a closed bijective realisation of $\widetilde{A}_{0}$.
The investigation and classification of mutually adjoint pairs $(A,A^{*})$ of closed bijective realisations of a Friedrichs operator $A_{0}$ (and of its partner $\widetilde{A}_{0}$) has been the object of multiple recent studies [8, 29, 30, 31, 32]. Quite noticeably ([8, Theorem 13]):

### v2

@@ 1. Introduction
This clearly encompasses a huge variety of concrete classes of inverse problems, including those where $A$ is a differential operator. On the one hand, it is customary to study each such class developing ad hoc methods and tools. This falls within the realm of numerical analysis, and includes the issue of reproducing the ‘exact’ solution(s) $f$ to $Af=g$ as a limit of approximants $f_{n}\to f$ obtained by solving (easier) truncated problems $A_{n}f_{n}=g_{n}$, typically in an iterative way.
@@
Obviously, there is a natural interplay between the two perspectives above. For instance, the actual convergence $f_{n}\to f$ from solutions to suitably truncated inverse problems obtained in a concrete setting, as well as the convergence rate and the quantitative control (say, the norm) of the displacement $f-f_{n}$, are often established by means of abstract functional-analytic and operator theoretic methods. We mention, in this respect, excellent classical references such as [44, 49, 55, 39, 27, 23, 40, 45, 51, 32, 8, 53, 50], and also, to quickly grasp the flavour of the abstract point of view, [17, Sections 1.1 and 1.2, and Appendix A].
@@
As a matter of fact, the theoretical apparatus developed so far in the subject of Krylov solvability is sufficiently articulated to make it natural to specialise it for distinguished classes of inverse problems. The one we consider here is the type of inverse problems customarily referred to as Friedrichs systems.
@@
Starting from concrete, differential formulations of Friedrichs-type inverse problems, it is possible to re-formulate them in an abstract Hilbert space setting, so as to exploit general operator-theoretic methods and then apply the results to concrete versions [32, 36, 1, 3, 7, 28, 29]. Section 3 revisits the basic aspects of this construction.
@@ 2.1. Krylov approximation methods as truncation algorithms
(2.3) $\mathcal{K}_{n}(A,g)\,:=\,\mathrm{span}\big{\{}g,Ag,A^{2}g,\dots,A^{n-1}g\big{\}}\,,\qquad\dim\mathcal{K}_{n}(A,g)\leqslant n\,,$
@@ 2.2. A glance at Krylov solvability for bounded inverse problems
Here Krylov reducibility indicates that $A$ is reduced (in the standard operator sense, e.g., [52, Section 1.4]) into two blocks with respect to the Hilbert space direct sum decomposition
@@
An evident sufficient condition for Krylov reducibility is self-adjointness: a bounded self-adjoint operator $A$ is always $\mathcal{K}(A,g)$-reduced for any $g\in\mathcal{H}$. More generally, bounded normal operators $A$ are $\mathcal{K}(A,g)$-reduced if and only if $A^{*}g\in\overline{\mathcal{K}(A,g)}$ [20, Proposition 2.1]. Thus, for unitary operators Krylov reducibility is equivalent to Krylov solvability [20, Remark 3.6].
@@
[20, Proposition 3.9] Let $A\in\mathcal{B}(\mathcal{H})$ and let $g\in\mathrm{ran}A$. If A is $\mathcal{K}(A,g)$-reduced, then there exists a Krylov solution to the problem $Af=g$. For example, if $f\in\mathcal{H}$ satisfies $Af=g$ and $P_{\mathcal{K}}$ is the orthogonal projection onto $\overline{\mathcal{K}(A,g)}$, then $f^{\circ}:=P_{\mathcal{K}}f$ is a Krylov solution
@@
[20, Proposition 3.4], [16, Proposition 2.4] Let $A\in\mathcal{B}(\mathcal{H})$ and let $g\in\mathrm{ran}A$. If $\mathcal{I}(A,g)=\{0\}$, then there exists a Krylov solution $f\in\overline{\mathcal{K}(A,g)}$ to the inverse linear problem $Af=g$. If, additionally, $A^{-1}\in\mathcal{B}(\mathcal{H})$, then $\mathcal{I}(A,g)=\{0\}$ if and only if the solution $f$ to $Af=g$ is in $\overline{\mathcal{K}(A,g)}$.
Observe that, obviously, the $\mathcal{K}(A,g)$-reducibility of an operator $A$ is just a special case of the triviality of the Krylov intersection $\mathcal{I}(A,g)$, which indicates that the latter subspace is more informative concerning the possibility that an inverse problem be Krylov-solvable.
@@
[22, Propositions 3.4 and 3.9] Let $A\in\mathcal{B}(\mathcal{H})$ be a compact normal operator and let $g\in\mathcal{H}$. Then $A$ is $\mathcal{K}(A,g)$-reduced. If, additionally, $g\in\mathrm{ran}A$, then $Af=g$ has a unique Krylov solution that is also a minimal norm solution to the inverse linear problem.
@@ 2.3. Krylov solvability and unbounded self-adjoint operators
and a discussion of its properties and relevance for the Krylov solvability of $Af=g$ may be found in [16, Sections 5 and 6].
@@
The self-adjoint case exhibits in this respect distinguished properties. For one thing, as well known (see, e.g., [52, Lemma 7.13]), if $A$ is self-adjoint, then $\mathcal{D}^{b}(A)$ is dense in $\mathcal{H}$, and so too are, consequently, all the other classes above. Moreover, a noticeable conclusion can be drawn concerning the Krylov solvability of possibly unbounded self-adjoint inverse problems.
@@
[16, Theorems 4.1 and 4.2] Let $A$ be a self-adjoint or skew-adjoint operator on a Hilbert space $\mathcal{H}$, and let $g\in\mathrm{ran}A\cap\mathcal{D}^{qa}(A)$. Then there exists a unique Krylov solution to the inverse linear problem $Af=g$.
It is worth commenting that Theorem 2.7 relies on an ingenious and somewhat laborious analysis carried out in the framework of the convergence of the conjugate gradient algorithm on an infinite-dimensional Hilbert space. This was initially developed in the bounded case [47, 48] and recently re-designed in the unbounded setting [18, Theorem 2.4 and Corollary 2.5]. The full proof obtained in [18] in fact reveals that the sequence of approximants from the conjugate-gradients algorithm converges to the minimal norm solution of the inverse problem (2.1) when $A$ is positive self-adjoint and $g\in\mathcal{D}^{qa}(A)\cap\mathrm{ran}A$. In other words: the truncation scheme and truncated inverse problems (2.4) defined by setting the trial space $Q_{n}=P_{n}$ for $P_{n}$ the orthogonal projection onto $\mathcal{K}_{n}(A,g)$, produce a well-defined sequence of approximants $(f_{n})_{n\in\mathbb{N}}$ and $f_{n}\xrightarrow{n\to\infty}f^{\circ}$ in $\mathcal{H}$, where $f^{\circ}$ is the minimal norm solution to (2.1).
@@ 2.4. Krylov subspaces and solvability under perturbations
[19, Example 3.3] Let $R:\ell^{2}(\mathbb{Z})\to\ell^{2}(\mathbb{Z})$ be the right shift operator $e_{n}\mapsto e_{n+1}$ (where $e_{n}$ is the $n$-th canonical vector of $\ell^{2}(\mathbb{Z})$). $R$ is unitary with $R^{*}=R^{-1}=L$, the left-shift operator $e_{n}\mapsto e_{n-1}$. $R$ admits a dense set $\mathcal{C}\subset\ell^{2}(\mathbb{Z})$ of cyclic vectors (i.e., vectors $g\in\mathcal{H}$ such that $\mathcal{K}(R,g)$ is dense in $\ell^{2}(\mathbb{Z})$) and a dense subset $\mathcal{N}\subset\ell^{2}(\mathbb{Z})$ of all finite linear combinations of the canonical vectors. Clearly all vectors in $\mathcal{N}$ are non-cyclic for the operator $R$, and moreover the solution $f$ to $Rf=g$ for any $g\in\mathcal{N}$ does not belong to $\overline{\mathcal{K}(R,g)}$.
@@
In a perturbative perspective, it is relevant to compare the possibly “close” Krylov subspaces $\mathcal{K}(A,g)$ and $\mathcal{K}(\widetilde{A},\widetilde{g})$. A promising tool is the weak gap metric between subspaces of separable Hilbert space, inspired to the standard norm gap distance (namely, the Hausdorff distance) between subspaces (see, e.g., [43, Chapt. 4, §2]). Owing to the separability, and also reflexivity, of $\mathcal{H}$, the Hilbert space weak topology restricted to the (norm-closed) unit ball $B_{\mathcal{H}}$ of $\mathcal{H}$ makes the ball a compact and complete metric space $(B_{\mathcal{H}},\rho_{w})$, and the metric $\rho_{w}$ is induced by a norm, namely $\rho_{w}(x,y)=\|x-y\|_{w}$ for some explicitly constructed norm in $B_{\mathcal{H}}$. This allows to define, for any pair of norm-closed linear subspaces $U$ and $V$ of $\mathcal{H}$,
@@
where $B_{U}=B_{\mathcal{H}}\cap U$ and $B_{V}=B_{\mathcal{H}}\cap V$. As a matter of fact [19, Lemmas 6.1 and 6.2], $\widehat{d}_{w}$ is indeed a metric for the norm-closed subspaces of $\mathcal{H}$, albeit in general it is not complete.
@@
of the $N$-dimensional truncated Krylov subspace $\mathcal{K}_{N}(A,g)$ to the (closure of the) full Krylov subspace $\mathcal{K}(A,g)$. Such quite natural inner Krylov approximability is actually false when the standard Hausdorff distance is used for an infinite-dimensional $\mathcal{H}$ [19, Lemma 7.1]. Another natural aspect of the inner Krylov approximability that manifests in the $\widehat{d}_{w}$-sense is the following [19, Lemma 7.6]:
@@
Other natural approximations, instead, do not hold in general in the $\widehat{d}_{w}$-sense, which suggests that further conditions need be added to the perturbation picture. For example [19, Lemma 7.3, Examples 7.4 and 7.5],
@@
[19, Proposition 7.7] Let $\mathcal{H}$ be a separable Hilbert space, and let the following be given:
@@ 3. Overview on abstract Friedrichs systems
They are currently being investigated in a variety of directions, including well-posedness [36, 1, 3, 7], homogenisation [12, 11], representations of boundary conditions [1], connection with the classical theory [2, 3, 6, 4], applications to various initial or boundary value problems of elliptic, hyperbolic, and parabolic type [5, 13, 25, 33, 46], development of different numerical schemes [10, 9, 14, 33, 34, 35], classification of abstract realisations of Friedrichs systems [7, 28, 29, 30, 31], the Friedrichs system structure of contact interactions [28], and many others.
@@
Many other concrete differential problems may be cast in the form $Af=g$ with $A$ of the type (3.1)-(3.2) [1, 2, 6, 5, 13, 33, 34]. In addition, it is possible to reformulate all such concrete problems at an abstract Hilbert space setting: this includes the classification of the distinct operator realisations of $A$ (namely, the counterpart of boundary conditions at $\partial\Omega$), as well as the identification of certain intrinsic abstract conditions ensuring the bijectivity of $A$ and hence the solvability of $Af=g$ [36, 1, 3, 7, 28, 29, 30, 31].
@@
In the abstract setting (3.3) it can be shown ([36, Section 2.1], [3, Theorem 7]) that the graph norms associated with $A_{0}$ and $\widetilde{A}_{0}$ are equivalent, and as a consequence $A_{0}$ and $\widetilde{A}_{0}$ have a common (dense) domain
@@
(3.6) $A_{1}+\widetilde{A}_{1}\,\subset\,\overline{A_{0}+\widetilde{A}_{0}}\,.$
@@
In particular, a special status is attributed to those $V$’s that are closed in $(\mathcal{W}_{1},\|\cdot\|_{A_{1}})$ and correspondingly those realisations $A_{1}|_{V}$’s that are bijective as $(V,\|\cdot\|_{A_{1}})\to\mathcal{H}$ maps. Indeed, such operators that are closed and densely defined in $\mathcal{H}$, are also $(V,\|\cdot\|_{A_{1}})\to\mathcal{H}$ bounded, and their inverse is $\mathcal{H}\to(V,\|\cdot\|_{A_{1}})$ bounded, implying that the inverse linear problem induced by each such $A_{1}|_{V}$ is well-posed. Incidentally, one can see that if $A:=A_{1}|_{V}$ is a closed bijective realisation of $A_{0}$, then its adjoint $A^{*}=(A_{1}|_{V})^{*}\equiv\widetilde{A}_{1}|_{\widetilde{V}}$, with $\widetilde{V}:=\mathrm{dom}\,(A_{1}|_{V})^{*}$, is a closed bijective realisation of $\widetilde{A}_{0}$.
The investigation and classification of mutually adjoint pairs $(A,A^{*})$ of closed bijective realisations of a Friedrichs operator $A_{0}$ (and of its partner $\widetilde{A}_{0}$) has been the object of multiple recent studies [7, 28, 29, 30, 31]. Quite noticeably ([7, Theorem 13]):